OpenStudy (anonymous):

I am with a group of five of my friends. A box contains 8 pieces each of milk chocolate, white chocolate, and dark chocolate. The box is passed around the six of us, with each person taking 4 pieces. Assume that each person chooses at random without replacement from the available pieces. I am the last person to whom the box is passed. Find the chance that I pick 4 dark chocolates.

OpenStudy (goformit100):

To solve these type of questions take Probability in use.

OpenStudy (anonymous):

8C4/24C4=(70/10626) 70/10626=

OpenStudy (anonymous):

thats wrong